OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 8? A. (0, 8), (1, 7), and (2, 9) B. (1, 8), (2, 16), and (3, 24) C. (1, 9), (3, 11), and (5, 13) D. (8, 0), (9, 1), and (10, 2)

OpenStudy (anonymous):

@Nameless hey can you hellp me with this one plz

OpenStudy (anonymous):

and this lol How many solutions are there to the equation y = 14x – 9? Explain your answer. A. There are no solutions because no value can be used for the variable x for the equation to remain true. B. There is one solution because only one value can be used for the variable x for the equation to remain true. C. There are 2 solutions because only 2 values can be used for the variables for the equation to remain true. D. There are an infinite number of solutions because any value can be used for the variable x and the equation would remain true.

OpenStudy (anonymous):

ok these are managable. the first one is like the last one we did. except you just add ... so its easier.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this time.. the x and y coordinates are different... like way different. so we need to look at each coordinate pair and quickly evaluate in our head whether or not the x value satisfies the equation.

OpenStudy (anonymous):

okay i understand that

OpenStudy (anonymous):

so y = x + 8. the first ordered pair (0, 8) when x = 0, does y = 8 (these should be able to be done rapidly in your head.)

OpenStudy (anonymous):

yes move to the next one if x = 1 does y = 7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So.. A.?

OpenStudy (anonymous):

substitute x and y ... y = 7, and x = 1... so 7=1+8 is true or false

OpenStudy (anonymous):

false

OpenStudy (anonymous):

so its not A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

substitute again ( 1, 8) does 8 = 1+8

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not B then. lets move onto C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Its D.

OpenStudy (anonymous):

y = x +8... Y is always greater than X .. right?

OpenStudy (anonymous):

Correct or Incorrect?

OpenStudy (anonymous):

well for positive integers of x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Does 1 = 9+8

OpenStudy (anonymous):

does 2 = 10+8

OpenStudy (anonymous):

it is not D.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so then its C.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

for the next one. are you aware that y=mx+b is the standard equation for a line.. where m is the slope (rise over run) of the line?

OpenStudy (anonymous):

No I wasnt aware of that

OpenStudy (anonymous):

so y=14x-9 is a straight line on a graph.

OpenStudy (anonymous):



OpenStudy (anonymous):

it looks like that

OpenStudy (anonymous):

a line goes in either direction until infinity right?

OpenStudy (anonymous):

and we saw from doing those past problems.. the points on the line (x, y) were simply the x and y in the equation.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

this is where it gets hard to follow. if you substitute in any value for x.. you will always get a value for y.. and that answer is going to represent a point on that line.

OpenStudy (anonymous):

okay i get that

OpenStudy (anonymous):

no matter what value you pick.. 1million.. 10 trillion... - 135quadrillion... or whatever number... even fractions.

OpenStudy (anonymous):

so lets look at the answers they give.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

based on your great knowledge in math that you now have, what do you guess would be the answer.

OpenStudy (anonymous):

These concepts are hard to understand and are by no means trivial. it took mathematicians years to figure this stuff out and you are expected to learn a book in a year. so no worries. it will come

OpenStudy (anonymous):

ok and I think B.

OpenStudy (anonymous):

y = 14x -9. if x = 0, y = 14*(0) -0 or y = 0 - 9 which is -9. that is true. if x = 1, y = 14 -9 which = 5. that is also true.. so it cannot be B. think about it. remember that I said you can put any number in for x.. 1/2 1/4 2,000 15,000 whatever number you want and it will be y... because y = whatever 14x-9 is equal to.

OpenStudy (anonymous):

and we know a line goes to infinity in either direction.

OpenStudy (anonymous):

oh ok ok

OpenStudy (anonymous):

y is 14x-9

OpenStudy (anonymous):

and x can be any number that you have learned about

OpenStudy (anonymous):

im thinking D. or C. now

OpenStudy (anonymous):

y = 14x -9. if x = 0, y = 14*(0) -0 or y = 0 - 9 which is -9. that is true. if x = 1, y = 14 -9 which = 5. that is also true.. if x = -1, y = 14(-1)-9 or y=-14-9 or y = -14 + (-9) which = negative 23. so now we have 3 solutions that we have used to make the equation true. as you can see.. any number you tell me we can put in to the equation for x whatever that is multiplied by 14 then - 9 is what y will be. you say.. okay.. i want x to be 10. okay lets let x = 10.. 14(10) -9 = 131..so y = 131. you can put in any number from negative infinity to positive infinity and this equation will be true... because of the fact that Y = whatever 14(x)-9 is equal to

OpenStudy (anonymous):

infinite solutions: means that there is no limit to the number of solutions

OpenStudy (anonymous):

if there is only one answer to the problem.. then the answer would be referred to as a finite solution.

OpenStudy (anonymous):

I hope that helps. I have to run for awhile. I would go with D.

OpenStudy (anonymous):

ok